The length of a rectangle is 2 times longer than its width. If the diagonal of the rectangle measures 21 inches, what is its length? Type your answer as a decimal rounded to the nearest hundredth.
Answer by amarjeeth123(569)

Let the width of the rectangle be x.
Then the length of the rectangle is 2x.
Diagonal of the rectangle =21 inches
Applying Pythagoras theorem,
(Diagonal)^2=(length)^2+(width)^2
x^2+(2x)^2=21^2
x^2+4x^2=441
5x^2=441
x^2=441/5=88.2
x=9.391485505499117
length=2x=18.78297101099823
The length of the rectangle is 18.783 inches